# Asymptotic condition numbers for linear ODEs

### v1

@@ 2.4.2. The matrix $zA$
The columns of the matrix $V(zA)$ corresponding to $zA$ are given in (20) and the rows of the matrix $W(zA)$ corresponding to $zA$ are given in (21). Now, use the formula (29). ∎
2.4.3. The real case
@@
Assume $A\in\mathbb{R}^{n\times n}$. Moreover, assume to have a Jordan basis as that described in Proposition 3.
@@ 2.4.4. The kernel of the matrices $P_{il}$
In other words, with reference to the previous Figure 2, the proposition says that the kernel of $P_{il}$ is constituted by the vectors $u\in\mathbb{C}^{n}$ with zero component along all the generalized eigenvectors $v^{(i,j,k)}$ whose third index $k$ (in the three-index notation) is an index column of elements in the intersection of $l$-th diagonal with mini-blocks.
@@
For $u\in\mathbb{C}^{n}$, we have
$P_{il}u=0$$\Leftrightarrow$ $V^{\left(i\right)}N^{(i,l)}\alpha^{\left(i\right)}\left(u\right)=0$
$\Leftrightarrow$ $N^{(i,l)}\alpha^{\left(i\right)}\left(u\right)=0$
by the linear independence of the vectors $v^{(i,j,k)}$ in (8)
$\Leftrightarrow$ $N^{(i,l,j)}\alpha^{\left(i,j\right)}\left(u\right)=0\ \text{for\ all\ }j=1,\ldots,d_{i},$
@@
 and for all $\text{\ and\ for all\ }k\in\{l+1,\ldots,m_{ij}\}.$
@@
The Proposition 13 confirms our previous observation in Subsection 2.4 that $P_{i0}$ is the projection on the generalized eigenspace corresponding to $\lambda_{i}$.
2.5. The index $l_{i}(u)$
In view of the Propositions 12 and 13, for $i\in\{1,\ldots,p\}$ and $u\in\mathbb{C}^{n}$ such that $\alpha^{(i)}(u)\neq 0$, we define the index
$l_{i}\left(u\right):=\max\left\{l\in\left\{0,\ldots,m_{i}-1\right\}:u\notin\mathrm{ker}\left(P_{il}\right)\right\}$
@@ 2.5. The index $l_{i}(u)$
by Propositions 12 and 13. In this case, the index $l_{i}(u)$ cannot be defined. On the other hand, if $\alpha^{(i)}(u)\neq 0$, then
@@
$u\notin\mathrm{ker}(P_{il}),\ l\in\{0,\ldots,l_{i}(u)\},$
@@
$u\in\mathrm{ker}(P_{il}),\ l\in\{l_{i}(u)+1,\ldots,m_{i}\}.$
Observe that if $\lambda_{i}$ is a non-defective eigenvalue, i.e. $m_{i}=0$, then $l_{i}\left(u\right)=0$.
@@
The following proposition gives $l_{i}\left(u\right)$ in terms of the zero components of $\alpha^{\left(i\right)}\left(u\right)$.
@@
In other words, the proposition says that $l_{i}(u)+1$ is the maximum index $k$, in a Jordan chain
@@
corresponding to the eigenvalue $\lambda_{i}$, such that $u$ has a non-zero component along the generalized eigenvector $v^{(i,j,k)}$. Alternatively, with reference to Figure 2, the proposition says that $l_{i}(u)$ is the maximum index $l$ such that $u$ has a non-zero component along some generalized eigenvector $v^{(i,j,k)}$ whose third index $k$ is an index column of elements in the intersection of $l$-th diagonal with mini-blocks.
@@
The proof is straightforward by recalling the interpretation of the statement of Proposition 14 from a few lines above, referencing Figure 2, and the observation about $\mathrm{ker}(P_{il})$, also referencing Figure 2, immediately following Proposition 11. However, a more formal proof is now presented below.
@@
Suppose $l_{i}\left(u\right)<m_{i}-1$ and then $u\notin\mathrm{ker}\left(P_{il_{i}(u)}\right)$ and $u\in\mathrm{ker}\left(P_{i\,l_{i}(u)+1}\right)$. By Proposition 11 and $u\notin\mathrm{ker}\left(P_{il_{i}(u)}\right)$, there exist $\overline{j}\in\{1,\ldots,d_{i}\}$ with $m_{i\overline{j}}\geq l_{i}(u)+1$ and $\overline{k}\in\{l_{i}\left(u\right)+1,\ldots,m_{i\overline{j}}\}$ such that $\alpha_{i\overline{j}\ \overline{k}}\neq 0$. Then
@@
On the other hand, by Proposition 11 and $u\in\mathrm{ker}\left(P_{i\,l_{i}(u)+1}\right)$, we have $\alpha_{ijk}\left(u\right)=0$for all $j\in\{1,\ldots,d_{i}\}$ with $m_{ij}\geq l_{i}(u)+2$ and for all $k\in\{l_{i}\left(u\right)+2,\ldots,m_{ij}\}$. Then
@@ 2.6. Linear independence
where the second inequality follows by (23). Since $V$ has linearly independent columns, we obtain, for $i\in\{1,\ldots,p\}$ such that $\alpha^{(i)}(u)\neq 0$,
00 0 $=$ $\sum\limits_{l=0}^{l_{i}\left(u\right)}c_{il}N^{(i,l)}\alpha^{\left(i\right)}\left(u\right)=\sum\limits_{l=0}^{l_{i}\left(u\right)}c_{il}\mathrm{diag}\left(N^{(i,l,1)},\ldots,N^{(i,l,d_{i})}\right)\alpha^{\left(i\right)}\left(u\right)$
@@
and then, for $j\in\{1,\ldots,d_{i}\}$,
$\sum\limits_{l=0}^{l_{i}\left(u\right)}c_{il}N^{(i,l,j)}\alpha^{\left(i,j\right)}\left(u\right)=0.$
@@
$=\left[\begin{array}[]{ccccccccc}c_{i0}&c_{i1}&\cdot&c_{il_{i}\left(u\right)-1}&c_{il_{i}\left(u\right)}&|&0&\cdot&0\\ 0&c_{i0}&c_{i1}&\cdot&c_{il_{i}\left(u\right)-1}&|&c_{il_{i}\left(u\right)}&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&|&c_{il_{i}\left(u\right)-1}&\cdot&0\\ \cdot&\cdot&\cdot&c_{i0}&c_{i1}&|&\cdot&\cdot&c_{il_{i}\left(u\right)}\\ \cdot&\cdot&\cdot&\cdot&c_{i0}&|&c_{i1}&\cdot&c_{il_{i}\left(u\right)-1}\\ -&-&-&-&-&&-&-&-\\ \cdot&\cdot&\cdot&\cdot&\cdot&|&c_{i0}&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&|&\cdot&\cdot&c_{i1}\\ 0&\cdot&\cdot&\cdot&\cdot&|&\cdot&0&c_{i0}\end{array}\right]\left[\begin{array}[]{c}\alpha_{i\overline{j}1}\left(u\right)\\ \alpha_{i\overline{j}2}\left(u\right)\\ \\ \cdot\\ \alpha_{i\overline{j}l_{i}\left(u\right)}\left(u\right)\\ \alpha_{i\overline{j}l_{i}\left(u\right)+1}\left(u\right)\\ -\\ 0\\ \cdot\\ 0\end{array}\right]$
@@
Now, since $\alpha_{i\overline{j}\,l_{i}\left(u\right)+1}\left(u\right)\neq 0$, we obtain $c_{il}=0$ successively for $l=0,1,\ldots,l_{i}\left(u\right)$.
@@
Consider $u\in\mathbb{C}^{n}$ such that $\alpha_{ijk}(u)\neq 0$ for $i\in\{1,\ldots,p\}$, $j\in\{1,\ldots,d_{i}\}$ and $k\in\{1,\ldots,m_{ij}\}$. For example, one can take $u=Ve$, where $e=(1,\ldots,1)$, and then $\alpha(u)=e$. We have $l_{i}(u)=m_{i}-1$ for $i\in\{1,\ldots,p\}$. Since
@@
and $P_{il}u$, $i\in\{1,\ldots,p\}$ and $l\in\{0,\ldots,l_{i}(u)\}$, are linearly independent (see previous Proposition 15), we have $c_{il}=0$, $i\in\{1,\ldots,p\}$ and $l\in\{0,\ldots,m_{i}-1\}$. ∎
2.6.1. Linear combinations depending on $t$
In the present paper, we deal with linear combinations of some of the vectors $P_{il}u$ in (32) or some of the matrices $P_{il}$ in (36), whose coefficients depend on time $t$ and they are not all zero. For a fixed $t$, such a linear combination is not zero since the vectors and the matrices are linearly independent (recall the two previous Propositions 15 and 16).
@@ 2.6.1. Linear combinations depending on $t$
[…]
@@
In the paper, we apply this lemma to the case where $a_{1},\ldots,a_{K}$ are some of the linearly independent vectors $P_{il}u$ in (32), or some of the linearly independent matrices $P_{il}$ in (36), and $f_{1}(t),\ldots,f_{K}(t)$ are of the form $\mathrm{e}^{\sqrt{-1}\omega t}$, with $\sqrt{-1}$ the imaginary unit and $\omega\in\mathbb{R}$, and then $\|f(t)\|_{\infty}=1$.
3. Asymptotic forms
In the next Section 4, we analyze the asymptotic (long-time) behavior of the three condition numbers (3), (4) and (5), i.e. what happens to (3), (4) and (5) as $t$ approaches infinity (becomes large).
This analysis determines the asymptotic forms of $\mathrm{e}^{tA}\widehat{z}_{0}$ and $\mathrm{e}^{tA}\widehat{y}_{0}$ in (3), of $\mathrm{e}^{tA}$ and $\mathrm{e}^{tA}\widehat{y}_{0}$ in (4) and of $\mathrm{e}^{tA}$ and $\mathrm{e}^{-tA}$ in (5). Then, the asymptotic behavior of the condition numbers is described by substituting in (3), (4) and (5) these asymptotic forms and by defining asymptotic condition numbers the new expressions obtained by these substitutions.
All of this might seem straightforward, but from a mathematical perspective it is not a simple task. Specifically:
  • It is necessary to precisely define what we mean by asymptotic form and how to determine it. This is particularly complicated in the non-diagonalizable case, where generalized eigenvectors are involved.
  • We also want to quantify how dominant the asymptotic form is at finite times.
  • The core of our analysis is the qualitative study of the asymptotic condition numbers. Understanding them requires significant mathematical effort in case of complex eigenvalues.
  • If we are interested in defining asymptotic condition numbers for the problem (2), the asymptotic condition numbers derived from (4) and (5) as $t\rightarrow+\infty$ may not be appropriate, since they represent the asymptotic worst cases of (3) and (4), respectively. Instead, we are concerned with the worst asymptotic cases of (3) and (4) by varying $\widehat{z}_{0}$ and $y_{0}$. In other words: do “asymptotic” and “worst” commute?
In this section, we determine the asymptotic forms of $\mathrm{e}^{tA}$ and $\mathrm{e}^{tA}u$, where $u$ is a vector. In order to define them some preliminary notions and results need to be introduced.
3.1. Notations $\sim$ and $\approx$
In this subsection, we precise what we mean for asymptotic form.
Let $f(t)$ and $g(t)$ be scalar, vector or matrix functions of $t\in\mathbb{R}$. For $g$ such that $g(t)\neq 0$ for $t$ in a neighborhood of +$\infty$+ ∞, we write
$f\left(t\right)\sim g\left(t\right),\ t\rightarrow+\infty,$ (38)
when
[…]
We interpret (38) as indicating that $g$ is an asymptotic form of $f$.
when
Observe that (39) means that the relative error of $f$ with respect to $g$ at the time $t$ is not larger than $\epsilon$. The notation (39) serves to quantify how much dominant the asymptotic form $g$ of $f$ is at the finite time $t$.
Remark 18.
Note that, for vector or matrix functions $f(t)$ and $g(t)$,
$f\left(t\right)\sim g\left(t\right),\ t\rightarrow+\infty,$
implies
and, for $\epsilon\geq 0$,
implies
This follows by
3.2. Partition of the spectrum and formula for the matrix exponential
where the $q$ distinct real parts $r_{j}$, $j\in\{1,\ldots,q\}$, of the eigenvalues of $A$ satisfy $r_{1}>r_{2}>\cdots>r_{q}$.
Figure 5. Partition of the eigenvalues $\lambda_{1},\ldots,\lambda_{p}$ of $A$ by decreasing real part. The eigenvalues are marked by “x”.
Now, we recall the formula (26) for the matrix exponential $\mathrm{e}^{tA}$:
where $m_{i}$ is the ascent of the eigenvalue $\lambda_{i}$ (maximum dimension of the mini-blocks corresponding to $\lambda_{i}$ in the JCF of $A$) and the matrices $P_{il}$ are defined and studied in Section 2. By using the partition of the spectrum given above, we have
Now, by exchanging the two inner sums we can rewrite (40) as
where
$L_{j}:=\max_{\lambda_{i}\in\Lambda_{j}}m_{i}-1,\ j\in\{1,\ldots,q\},$
and
3.3. Notations for the asymptotic form of $\mathrm{e}^{tA}u$
For $j\in\{1,\ldots,q\}$ and $u\in\mathbb{C}^{n}$:
  • Let
    $\Lambda_{j}(u):=\{\lambda_{i}\in\Lambda_{j}:\alpha^{(i)}(u)\neq 0\},$
    where $\alpha^{(i)}(u)$, defined in Section 2, is the vector of the Jordan basis components of $u$ along the generalized eigenvectors corresponding to $\lambda_{i}$. In other words, $\Lambda_{j}(u)$ is obtained from $\Lambda_{j}$ by including only the eigenvalues for which $u$ has a non-zero projection onto the corresponding generalized eigenspace.
  • In the case where $\Lambda_{j}(u)\neq\emptyset$, let
    $L_{j}(u):=\max\{l_{i}(u):\lambda_{i}\in\Lambda_{j}(u)\},$
    where
    $l_{i}(u)=\max\{l\in\{0,\ldots,m_{i}-1\}:P_{il}u\neq 0\}$
    is defined and studied in Section 2, and it is such that $l_{i}(u)+1$ is the maximum index $k$, in a Jordan chain
    $(v^{(i,j,k)})_{k=1,\ldots,m{ij}},\ j\in\{1,\ldots,d_{j}\},$
    corresponding to the eigenvalue $\lambda_{i}$, for which $u$ has a non-zero component along the generalized eigenvector $v^{(i,j,k)}$. In other words, $L_{j}(u)+1$ is the maximum index $k$, in a Jordan chain corresponding to an eigenvalue in $\Lambda_{j}$, for which $u$ has a non-zero component along the $k$-th generalized eigenvector of the chain.
3.3.1. Indeces of dominance
For $u\in\mathbb{C}^{n}\setminus\{0\}$:
  • Let
    In other words, $r_{j(u)}$ is the rightmost real part of an eigenvalue of $A$ such that $u$ has a non-zero projection on the corresponding generalized eigenspace. We call $j(u)$ the primary index of dominance of $u$.
  • Let
    $L(u):=L_{j(u)}(u).$
    In other words, $L(u)+1$ is the maximum index $k$, in a Jordan chain corresponding to an eigenvalue in $\Lambda_{j(u)}(u)$, for which $u$ has a non-zero component along the $k$-th generalized eigenvector of the chain. We call $L(u)$ the secondary index of dominance of $u$.
Similarly to Section 2, the next two subsections, 3.4 and 3.5, can be skipped and referred back to when necessary.
3.4. Properties of the matrices $Q_{jl}(t)$

### v2

@@ A.4.2. The matrix zAzA
Recall Proposition 24. The columns of the matrix $V(zA)$ corresponding to $zA$ are given in (43) and the rows of the matrix $W(zA)$ corresponding to $zA$ are given in (44). Now, use the formula (52). ∎
A.4.3. The real case
@@
Assume $A\in\mathbb{R}^{n\times n}$. Moreover, assume we have a Jordan basis as that described in Proposition 22.
@@ A.4.4. The kernel of the matrices $P_{il}$
Remark 32.
With reference to the previous Figure 6 and 7, Proposition 31 states that the kernel of $P_{il}$ is constituted by the vectors $u\in\mathbb{C}^{n}$ with zero component $\alpha_{ijk}(u)$ along all the generalized eigenvectors $v^{(i,j,k)}$ such that, in the three-index notation, $(i,j,k)$ is an index column of elements in the intersection of the $l$-th diagonal of $N^{(i,l)}$ with mini-blocks.
@@
For $u\in\mathbb{C}^{n}$, by Proposition 27 we have
$P_{il}u=0$ $\Leftrightarrow$ $\sum\limits_{\begin{subarray}{c}j=1\\ m_{ij}\geq l+1\end{subarray}}^{d_{i}}\sum\limits_{k=1}^{m_{ij}-l}\alpha_{ij\,l+k}(u)v^{\left(i,j,k\right)}=0$
@@
and for all $k\in\{l+1,\ldots,m_{ij}\}$,
where the second $\Leftrightarrow$ follows by the linear independence of the vectors $v^{(i,j,k)}$ in (31). ∎
@@
Proposition 34 confirms our previous observation in Subsection A.3.1 that $P_{i0}$ is the projection on the generalized eigenspace corresponding to $\lambda_{i}$.
A.5. The index $l_{i}(u)$
In view of the Propositions 33 and 34, for $i\in\{1,\ldots,p\}$ and $u\in\mathbb{C}^{n}$ such that $\alpha^{(i)}(u)\neq 0$, we define the index
$l_{i}\left(u\right):=\max\left\{l\in\left\{0,\ldots,m_{i}-1\right\}:u\notin\mathrm{ker}\left(P_{il}\right)\right\}$ (54)
@@ A.5. The index $l_{i}(u)$
by Propositions 33 and 34. In this case, the index $l_{i}(u)$ cannot be defined. On the other hand, if $\alpha^{(i)}(u)\neq 0$, then
@@
$u\notin\mathrm{ker}(P_{il}),\ l\in\{0,\ldots,l_{i}(u)\}$
@@
$u\in\mathrm{ker}(P_{il}),\ l\in\{l_{i}(u)+1,\ldots,m_{i}-1\},$
as illustrated in Figure 8.
Observe that if $\lambda_{i}$ is a non-defective eigenvalue, i.e., $m_{i}=1$, and $\alpha^{(i)}(u)\neq 0$, then $l_{i}\left(u\right)=0$.
@@
Remark 35.
By recalling the definition (54) of $l_{i}(u)$ and Remark 32 regarding $\mathrm{ker}(P_{il})$ and referencing Figures 6 and 7, we can say that $l_{i}(u)$ is the maximum index $l$ for which $u$ has a non-zero component $\alpha_{ijk}(u)$ along some generalized eigenvector $v^{(i,j,k)}$ such that, in the three-index notation, $(i,j,k)$ is an index column of elements in the intersection of the $l$-th diagonal of $N^{(i,l)}$ with mini-blocks.
The following proposition relates $l_{i}(u)$ to Jordan chains.
@@
In other words, the proposition states that $l_{i}(u)+1$ is the maximum index $k$ such that $u$ has non-zero component $\alpha_{ijk}(u)$ along the $k$-th generalized eigenvector $v^{(i,j,k)}$ of some Jordan chain
@@
corresponding to the eigenvalue $\lambda_{i}$.
@@
The proof should be straightforward by recalling Remark 35. Nevertheless, a more formal proof is presented below.
@@
Suppose $l_{i}\left(u\right)<m_{i}-1$ and then $u\notin\mathrm{ker}\left(P_{il_{i}(u)}\right)$ and $u\in\mathrm{ker}\left(P_{i\,l_{i}(u)+1}\right)$. By Proposition 31 and $u\notin\mathrm{ker}\left(P_{il_{i}(u)}\right)$, there exist $\overline{j}\in\{1,\ldots,d_{i}\}$ with $m_{i\overline{j}}\geq l_{i}(u)+1$ and $\overline{k}\in\{l_{i}\left(u\right)+1,\ldots,m_{i\overline{j}}\}$ such that $\alpha_{i\overline{j}\ \overline{k}}(u)\neq 0$. Then
@@
On the other hand, by Proposition 31 and $u\in\mathrm{ker}\left(P_{i\,l_{i}(u)+1}\right)$, we have $\alpha_{ijk}\left(u\right)=0\$for all $j\in\{1,\ldots,d_{i}\}$ with $m_{ij}\geq l_{i}(u)+2$ and for all $k\in\{l_{i}\left(u\right)+2,\ldots,m_{ij}\}$. Then
@@ A.6. Linear independence
where the second equality follows by (46). Since $V$ has linearly independent columns, we obtain, for $i\in\{1,\ldots,p\}$ such that $\alpha^{(i)}(u)\neq 0$,
$0$ $=$ $\sum\limits_{l=0}^{l_{i}\left(u\right)}c_{il}N^{(i,l)}\alpha^{\left(i\right)}\left(u\right)=\sum\limits_{l=0}^{l_{i}\left(u\right)}c_{il}\mathrm{diag}\left(N^{(i,l,1)},\ldots,N^{(i,l,d_{i})}\right)\alpha^{\left(i\right)}\left(u\right)$
@@
and then
$\sum\limits_{l=0}^{l_{i}\left(u\right)}c_{il}N^{(i,l,j)}\alpha^{\left(i,j\right)}\left(u\right)=0,\ j\in\{1,\ldots,d_{i}\}.$
@@
${\footnotesize=\left[\begin{array}[]{ccccccccc}c_{i0}&c_{i1}&\cdot&c_{il_{i}\left(u\right)-1}&c_{il_{i}\left(u\right)}&|&0&\cdot&0\\ 0&c_{i0}&c_{i1}&\cdot&c_{il_{i}\left(u\right)-1}&|&c_{il_{i}\left(u\right)}&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&|&c_{il_{i}\left(u\right)-1}&\cdot&0\\ \cdot&\cdot&\cdot&c_{i0}&c_{i1}&|&\cdot&\cdot&c_{il_{i}\left(u\right)}\\ \cdot&\cdot&\cdot&\cdot&c_{i0}&|&c_{i1}&\cdot&c_{il_{i}\left(u\right)-1}\\ -&-&-&-&-&&-&-&-\\ \cdot&\cdot&\cdot&\cdot&\cdot&|&c_{i0}&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&|&\cdot&\cdot&c_{i1}\\ 0&\cdot&\cdot&\cdot&\cdot&|&\cdot&0&c_{i0}\end{array}\right]\left[\begin{array}[]{c}\alpha_{i\overline{j}1}\left(u\right)\\ \alpha_{i\overline{j}2}\left(u\right)\\ \\ \cdot\\ \alpha_{i\overline{j}l_{i}\left(u\right)}\left(u\right)\\ \alpha_{i\overline{j}l_{i}\left(u\right)+1}\left(u\right)\\ -\\ 0\\ \cdot\\ 0\end{array}\right]}$
@@
Since $\alpha_{i\overline{j}\,l_{i}\left(u\right)+1}\left(u\right)\neq 0$, we obtain $c_{il}=0$ successively for $l=0,1,\ldots,l_{i}\left(u\right)$.
@@
Consider $u\in\mathbb{C}^{n}$ such that $\alpha_{ijk}(u)\neq 0$ for $i\in\{1,\ldots,p\}$, $j\in\{1,\ldots,d_{i}\}$ and $k\in\{1,\ldots,m_{ij}\}$. For example, one can take $u=Ve$, where $e=(1,\ldots,1)$, and then $\alpha(u)=e$. We have $l_{i}(u)=m_{i}-1$ for $i\in\{1,\ldots,p\}$. Since we have
@@
and $P_{il}u$, $i\in\{1,\ldots,p\}$ and $l\in\{0,\ldots,l_{i}(u)\}$, are linearly independent by the previous Proposition 37, we obtain $c_{il}=0$, $i\in\{1,\ldots,p\}$ and $l\in\{0,\ldots,m_{i}-1\}$. ∎
A.6.1. Linear combinations depending on $t$
In the present paper, we deal with linear combinations of some of the vectors $P_{il}u$ in (56), or some of the matrices $P_{il}$ in (60), whose coefficients depend on time $t$ and they are not all zero. For a fixed $t$, such a linear combination is not zero since the vectors and the matrices are linearly independent: recall the two previous Propositions 37 and 38.
@@ A.6.1. Linear combinations depending on $t$
Let $\mathcal{V}$ be a vector space over $\mathbb{C}$ equipped with the norm $\left\|\ \cdot\ \right\|$, let $a_{1},\ldots,a_{K}\in\mathcal{V}$ and let $f:I\rightarrow\mathbb{C}^{K}$, where $I$ is an arbitrary set. For any $t\in I$, the components $f_{1}(t),\ldots,f_{K}(t)$ of $f(t)$ serve as coefficients in the linear combination
@@
In the present paper, we apply this lemma to the case where $a_{1},\ldots,a_{K}$ are some of the linearly independent vectors $P_{il}u$ in (56), or some of the linearly independent matrices $P_{il}$ in (60), and $f_{1}(t),\ldots,f_{K}(t)$ are of the form $\mathrm{e}^{\sqrt{-1}\omega t}$, with $\sqrt{-1}$ the imaginary unit and $\omega\in\mathbb{R}$, and then $\|f(t)\|_{\infty}=1$.
Appendix B Properties of the matrices $Q_{jl}(t)$